# On the Optimality of Coded Distributed Computing for Ring Networks

### v1

@@ Abstract
We consider a coded distributed computing problem in a ring-based communication network, where $\mathsf{N}$ computing nodes are arranged in a ring topology and each node can only communicate with its neighbors within a constant distance $d$. To mitigate the communication bottleneck in exchanging intermediate values, we propose new coded distributed computing schemes for the ring-based network that exploit both ring topology and redundant computation (i.e., each map function is computed by $r$ nodes). Two typical cases are considered: all-gather where each node requires all intermediate values mapped from all input files, and all-to-all where each node requires a distinct set of intermediate values from other nodes. For the all-gather case, we propose a new coded scheme based on successive reverse carpooling where nodes transmit every encoded packet containing two messages traveling in opposite directions along the same path. Theoretical converse proof shows that our scheme achieves the optimal tradeoff between communication load, computation load $r$, and broadcast distance $d$ when $\mathsf{N}\gg d$. For the all-to-all case, instead of simply repeating our all-gather scheme, we delicately deliver intermediate values based on their proximity to intended nodes to reduce unnecessary transmissions. We derive an information-theoretic lower bound on the optimal communication load and show that our scheme is asymptotically optimal under the cyclic placement when $\mathsf{N}\gg r$. The optimality results indicate that in ring-based networks, the redundant computation $r$ only leads to an additive gain in reducing communication load while the broadcast distance $d$ contributes to a multiplicative gain.
Index Terms:
Coded computing, distributed computation, ring networks, communication load.
@@ I Introduction
  • For all-to-all computing, we proposed a coded transmission scheme that achieves the NCL of $O(\frac{(\mathsf{N}-r)^{2}}{8d})$, and is asymptotically optimal under the cyclic file placement, [^1] The cyclic placement has a simple form to be implemented, and is independent of the computation tasks, making it suitable for offline scenarios. Therefore, it has been widely adopted in many coded distributed computing, including [26, 27, 28]. when $\mathsf{N}$ is relatively large compared to $r$. Furthermore, we propose another scheme that is optimal when $d=1$ and $r\geq\mathsf{N}/2$, regardless of file placement.
  • From the optimality results of both all-gather and all-to-all computing, an interesting insight is that the computation load $r$ only leads to an additive gain in the reduction of communication load, while the broadcast distance $d$ contributes to a multiplicative gain. This finding differs from prior results on coded distributed problems [9, 10, 11, 12, 13, 14, 15, 16, 17], where the computation load $r$ could lead to multiplicative gains.
@@
Notation: We use the sans serif font for constants, bold for vectors and matrices, and calligraphic font for most sets. Let $\mathbb{N}^{+}$ denote the set of positive integers. For any $k\in\mathbb{N}^{+}$, define $[k]\triangleq\{1,2,\ldots,k\}$. For any integer $a$, we define $a_{\bmod\mathsf{N}}\triangleq\tilde{a}\in[\mathsf{N}]$ such that $(a-\tilde{a})\bmod{\mathsf{N}}=0$. For a vector $\left[X^{(1)},\ldots,X^{(t)}\right]$, we denote it as $X^{t}$. The operation $\oplus$ denotes addition over a finite field.
@@ II-A System Model
The $\mathsf{N}$ nodes are assigned some computing tasks over $\mathsf{N}$ input files $w_{1},\ldots,w_{\mathsf{N}}\in\mathbb{F}_{q}^{\mathsf{F}}$ for some $\mathsf{F}\in\mathbb{N}^{+}$, where $\mathbb{F}_{q}$ denotes a finite field of order $q$. Each node $n_{k}$ initially caches a set of the files $\mathcal{M}_{k}\subseteq\{w_{1},\ldots,w_{\mathsf{N}}\}$. We assume that each input file is at least cached by one node, i.e., $\bigcup\limits_{k=1,\ldots,\mathsf{N}}\mathcal{M}_{k}=\{w_{1},\ldots,w_{\mathsf{N}}\}$. Node $n_{k}$ first locally computes the functions that map the files in the set $\mathcal{M}_{k}$ into some IVs, then exchanges IVs with neighboring nodes to compute its output function. We define the computation load as follows.
@@ II-A1 All-Gather
where $g_{1,i}$ is a map function $\mathbb{F}_{q}^{\mathsf{F}}\to\mathbb{F}_{q}^{\mathsf{B}_{1}}$ for file $w_{i}$. After clock tick $t-1$ and before clock tick $t$ for $t=1,\ldots,\mathsf{T_{\text{e}}}$, node $n_{k}$ creates a coding symbol $X_{k}^{(t)}\in\mathbb{F}_{q}^{l_{k}^{(t)}}$, for some $l_{k}^{(t)}\in\mathbb{N}$, as a function of IVs computed locally and the received messages from pastime clock $Y_{k}^{(t-1)},\ldots,Y_{k}^{(1)}$, i.e.,
@@ II-A2 All-to-All
We are interested in designing efficient transmission schemes for these two problems. We introduce the following metric to formulate the communication load.
@@
Given the computation load $r$ and broadcast distance $d$, we define the normalized communication loads for the the all-gather and all-to-all computing problems as $\mathsf{T}_{1}(r,d)=\frac{\sum_{t\in[\mathsf{T_{\text{e}}}]}\sum_{k\in[\mathsf{N}]}l_{k}^{(t)}}{\mathsf{N}\mathsf{B}_{1}}$ and $\mathsf{T}_{2}(r,d)=\frac{\sum_{t\in[\mathsf{T_{\text{e}}}]}\sum_{k\in[\mathsf{N}]}l_{k}^{(t)}}{\mathsf{N}\mathsf{B}_{2}}$, respectively. That is, the NCL represents the normalized number of bits communicated by each node in the network. Define the minimum NCLs for the all-gather and all-to-all computing problems as $\mathsf{T}_{1}^{\star}(r,d)$ and $\mathsf{T}_{2}^{\star}(r,d)$, respectively. The objective of this paper is to characterize $\mathsf{T}_{1}^{\star}(r,d)$ and $\mathsf{T}_{2}^{\star}(r,d)$, given the computation load $r$ and broadcast distance $d$.
@@ II-B1 Reverse Carpooling
Our transmission strategy is built on reverse carpooling [30] that realizes the benefits of network coding. As demonstrated in Fig. 3 (a), where nodes $n_{1}$ and $n_{3}$ want to exchange packets through a relay node $n_{2}$. After receiving packets $P_{a}$ and $P_{b}$ sent from two different nodes on the opposite sides, the relay $n_{2}$ broadcasts the linear combination $P_{1}\oplus P_{2}$. Each of $n_{1}$ and $n_{3}$ can then obtain the desired packet by subtracting the packet it sent previously from the mixed packet. Therefore, with network coding, it needs $3$ transmissions. In contrast, the traditional forwarding approach requires $4$ transmissions (two for each packet). In more general, reverse carpooling involves two information flows that traverse a path in opposite directions. For example, as shown in Fig. 3 (b), a common path $(n_{1},n_{2},n_{3},n_{4})$ is shared by two flows, from $n_{1}$ to $n_{4}$ and from $n_{4}$ to $n_{1}$. It can be verified that the reverse carpooling technique can save up to $50\%$ of transmission. There will be multiple information flows in our ring network. The reverse carpooling technique can benefit the transmission when some broadcasts form the structure as Fig. 3(a).
@@
Figure 3: (a) Reverse carpooling with $3$ nodes. (b) Reverse carpooling for two flows. $P_{a}^{(t)}$ and $P_{b}^{(t)}$ are the packets sent by nodes $n_{1}$ and $n_{4}$ at time clock $t$, respectively. At clock $t$, node $n_{2}$ broadcasts $P_{a}^{(t-1)}\oplus P_{b}^{(t-2)}$, and node $n_{3}$ broadcasts $P_{a}^{(t-2)}\oplus P_{b}^{(t-1)}$. It effectively enables the two flows to traverse a common path without interfering with each other.
@@ II-B2 Efficient Broadcasting Using Network Coding
Following Algorithm 1, at the end of round $k$, each node in $A$ and $B$ can receive two new symbols from the two source nodes that are $2k$ and $12k-$ 1 nodes away along the circle, respectively. The algorithm will finish after at most $\frac{\mathsf{N}}{4}$ rounds, and all nodes can decode each input symbol. We can conclude that the NCL is
@@
    if $k>1$, each $\alpha_{i}\in A$ transmits the sum of the two information symbols it received in phase $2$, round $k-1$.
@@ III Main Results
For the all-gather computing system with a computation load $r\in\{1,\ldots,\mathsf{N}\}$ and a broadcast distance $d\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$ over the considered ring networks with $\mathsf{N}$ nodes, the following NCL is achivable
@@
$\mathsf{T}_{1}^{\star}(r,d)\geq\frac{\mathsf{N}-r}{2d}.$ (7)
Moreover, we have
@@
The lower convex of all points $\big{\{}\big{(}r,\left\lceil\frac{\mathsf{N}-r}{2d}\right\rceil\big{)}:r\in\{1,\ldots,\mathsf{N}\}\big{\}}$ is achievable for general $1\leq r\leq\mathsf{N}$ by using memory-sharing. Meanwhile, the lower bound in (7) holds for arbitrary file placement.
@@
For the all-to-all computing over the ring-based networks, we obtain the achievable NCL in Theorem 2 and a lower bound on NCL for $r<\frac{\mathsf{N}}{2}$ under the cyclic placement in Theorem 3.
@@
$\mathsf{T}_{\textnormal{2-cyc}}^{\textnormal{ach}}(r,d)=\left\{\begin{array}[]{cc}\vspace{5pt}O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{3(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8d}\right)&\!2r-1\leq\!d\!\\ \vspace{5pt}\!\!\!O\!\!\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!3r)\!+\!\frac{3\mathsf{N}-5r}{4}\!+\!\frac{9r^{2}+4(\mathsf{N}-r+1)}{8d}\right)&\!\!r\!-\!1\!<\!d\!\leq\!2(r\!-\!1)\\ O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8d}\right)&\!d\!\leq r\!-\!1\end{array}\right.$ (12)
@@
For the all-to-all computing system with a given computation load $r\in\{1,\ldots,\lceil\frac{\mathsf{N}}{2}\rceil-1\}$ and a broadcast distance $d\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$ over the considered ring networks with $\mathsf{N}$ nodes, a lower bound for the NCL under cyclic placement is given as:
$\mathsf{T}_{\textnormal{2-cyc}}^{\star}(r,d)\geq\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}-r+1).$ (13)
The scheme in Theorem 2 is asymptotically optimal when $\mathsf{N}\gg r$.
@@
Theorem 3 presents the converse result for the case $r\leq\lceil\frac{\mathsf{N}}{2}\rceil-1$ under the cyclic placement. When the computation load is relatively large $r\in\{\lceil\frac{\mathsf{N}}{2}\rceil,\ldots,\mathsf{N}\}$, we present the following optimal result when $d=1$, without any constraints on data placement.

### v2

@@ Abstract
We consider a coded distributed computing problem in a ring-based communication network, where $\mathsf{N}$ computing nodes are arranged in a ring topology and each node can only communicate with its neighbors within a constant distance $d$. To mitigate the communication bottleneck in exchanging intermediate values, we propose new coded distributed computing schemes for the ring-based network that exploit both ring topology and redundant computation (i.e., each map function is computed by $r$ nodes). Two typical cases are considered: all-gather where each node requires all intermediate values mapped from all input files, and all-to-all where each node requires a distinct set of intermediate values from other nodes. For the all-gather case, we propose a new coded scheme based on successive reverse carpooling, where nodes transmit every encoded packet containing two messages traveling in opposite directions along the same path. Theoretical converse proof shows that our scheme achieves the optimal tradeoff between communication load, computation load $r$, and broadcast distance $d$ when $\mathsf{N}\gg d$. For the all-to-all case, instead of simply repeating our all-gather scheme, we delicately deliver intermediate values based on their proximity to intended nodes to reduce unnecessary transmissions. We derive an information-theoretic lower bound for arbitrary file placements and show that our scheme is asymptotically optimal under the cyclic placement when $\mathsf{N}\gg r$. The optimality results indicate that in ring-based networks, the redundant computation $r$ only leads to an additive gain in reducing communication load while the broadcast distance $d$ contributes to a multiplicative gain.
@@ I Introduction
  • For all-to-all computing, we proposed a coded transmission scheme that achieves the NCL of $O(\frac{(\mathsf{N}-r)^{2}}{8d})$, and is asymptotically optimal under the cyclic file placement, [^1] The cyclic placement has a simple form to be implemented, and is independent of the computation tasks, making it suitable for offline scenarios. Therefore, it has been widely adopted in many coded distributed computing, including [26, 27, 28]. when $\mathsf{N}$ is relatively large compared to $r$. We also propose another scheme that is optimal when $d=1$ and $r\geq\mathsf{N}/2$, regardless of file placement. In addition, we establish a placement-agnostic converse bound that provides a general performance benchmark for arbitrary file assignments.
  • From the optimality results of both all-gather and all-to-all computing, an interesting insight is that the computation load $r$ only leads to an additive gain in the reduction of communication load, while the broadcast distance $d$ contributes to a multiplicative gain. This behavior differs from prior results on coded distributed problems [9, 10, 11, 12, 13, 14, 15, 16, 17], where the computation load $r$ could lead to multiplicative gains.
Many recent works in coded caching [29, 30], coded distributed computing [9, 31], and linearly separable computing [32] have demonstrated communication-computation gains under arbitrary, pre-set, or decentralized placements, typically assuming a shared link or fully connected network. Topology-aware coded distributed computing has further incorporated network constraints into the system design, such as fat-tree architectures [16]. In comparison, the present work focuses on ring networks with limited broadcast distance $d$, where multicast opportunities are structurally restricted by topology, and the cyclic placement is adopted as a topology-aligned design. Notably, when $d$ is sufficiently large ($d\geq\lfloor\mathsf{N}/2\rfloor$), the ring network turns to a fully connected topology; for smaller $d$, the stronger topological constraints render existing shared-link schemes inapplicable without substantial modification.
@@
Notation: We use sans-serif font for constants, boldface for vectors and matrices, and calligraphic font for sets. Let $\mathbb{N}^{+}$ denote the set of positive integers. For any $k\in\mathbb{N}^{+}$, define $k]\triangleq\{1,2,\ldots,k\}$. For any integer $a$, define $a_{\bmod\mathsf{N}}\triangleq\tilde{a}\in[\mathsf{N}]$ such that $(a-\tilde{a})\bmod\mathsf{N}=0$. For a vector $X^{(1)},\ldots,X^{(t)}]$, we write $X^{t}$ for brevity. The operation $\oplus$ denotes addition over a finite field.
TABLE I: Summary of Main Notation
Symbol Description
$\mathsf{N}$ Number of computing nodes in the ring network
$n_{i}$ Node $i$, $i\in[\mathsf{N}]$
$w_{j}$ Input file $j$, $j\in[\mathsf{M}]$
$r$ Computation load (each file mapped by $r$ nodes on average)
$d$ Broadcast distance in the ring topology
$V_{i}$ Intermediate value (IV) of file $i$ needed by all nodes in all-gather problem
$v_{i}^{j}$ Intermediate value (IV) of file $j$ needed by node $i$ in all-to-all problem
$\mathsf{T}_{1}(r,d)$ NCL for all-gather computing
$\mathsf{T}_{2}(r,d)$ NCL for all-to-all computing
$X_{i}^{(t)}$ Coded transmission from node $n_{i}$ at clock tick $t$
$X_{i}^{(j,k)}$ Coded transmission from node $n_{i}$ in round $j$, step $k$
@@ II-A System Model
The $\mathsf{N}$ nodes are assigned some computing tasks over $\mathsf{N}$ input files $w_{1},\ldots,w_{\mathsf{N}}\in\mathbb{F}_{q}^{\mathsf{F}}$ for some $\mathsf{F}\in\mathbb{N}^{+}$, where $\mathbb{F}_{q}$ denotes a finite field of order qq. [^3] A system with an arbitrary number of input files and output functions can be reduced to this setting via appropriate padding and batching. See Remark 3 for further discussion. Each node $n_{k}$ initially caches a set of the files $\mathcal{M}_{k}\subseteq\{w_{1},\ldots,w_{\mathsf{N}}\}$. We assume that each input file is at least cached by one node, i.e., $\bigcup\limits_{k=1,\ldots,\mathsf{N}}\mathcal{M}_{k}=\{w_{1},\ldots,w_{\mathsf{N}}\}$. Node $n_{k}$ first locally computes the functions that map the files in the set $\mathcal{M}_{k}$ into some IVs, then exchanges IVs with neighboring nodes to compute its output function. We define the computation load as follows.
@@ II-A1 All-Gather
where $g_{1,i}$ is a map function $\mathbb{F}_{q}^{\mathsf{F}}\to\mathbb{F}_{q}^{\mathsf{B}_{1}}$ for file $w_{i}$. After clock tick t1t-1 and before clock tick $t$ for $t=1,\ldots,\mathsf{T_{\text{e}}}$, node $n_{k}$ creates a coding symbol $X_{k}^{(t)}\in\mathbb{F}_{q}^{l_{k}^{(t)}}$, for some $l_{k}^{(t)}\in\mathbb{N}$, as a function of IVs computed locally and the received messages from pastime clock $Y_{k}^{(t-1)},\ldots,Y_{k}^{(1)}$, i.e.,
@@ II-A2 All-to-All
Consistent with the literature on coded distributed computing, we assume that data placement occurs in a preprocessing phase. Our focus is on designing efficient transmission schemes for the delivery of IVs, which typically constitutes the communication bottlenecks. To this end, we introduce the following metric to formulate the communication load.
@@
Given the computation load $r$ and broadcast distance $d$, we define the normalized communication loads for the all-gather and all-to-all computing problems as $\mathsf{T}_{1}(r,d)=\frac{\sum_{t\in[\mathsf{T_{\text{e}}}]}\sum_{k\in[\mathsf{N}]}l_{k}^{(t)}}{\mathsf{N}\mathsf{B}_{1}}$ and $\mathsf{T}_{2}(r,d)=\frac{\sum_{t\in[\mathsf{T_{\text{e}}}]}\sum_{k\in[\mathsf{N}]}l_{k}^{(t)}}{\mathsf{N}\mathsf{B}_{2}}$, respectively. That is, the NCL represents the normalized number of bits communicated by each node in the network. Define the minimum NCLs for the all-gather and all-to-all computing problems as $\mathsf{T}_{1}^{\star}(r,d)$ and $\mathsf{T}_{2}^{\star}(r,d)$, respectively. The objective of this paper is to characterize $\mathsf{T}_{1}^{\star}(r,d)$ and $\mathsf{T}_{2}^{\star}(r,d)$, given the computation load $r$ and broadcast distance $d$.
These two problems require fundamentally new transmission and decoding designs compared with existing coded computing frameworks, due to the distance-limited broadcast constraint imposed by the ring topology. Moreover, the all-to-all problem is inherently more challenging than all-gather because information flows are directional. As will be shown in the subsequent sections, addressing these challenges requires the integration of multi-round scheduling, topology-aware coding, and successive decoding mechanisms tailored to the ring structure.
@@
(Latency) Consider a full-parallelism model where all nodes transmit simultaneously over orthogonal resources (e.g., in time, frequency, or code domains) without mutual interference. The latency can be characterized by $L(r,d)=\frac{\sum_{t\in[\mathsf{T_{\text{e}}}]}\max\left(l_{1}^{(t)},\dots,l_{\mathsf{N}}^{(t)}\right)}{\mathsf{B}}$, where $\mathsf{B}$ denotes per-node transmission rate (in bits per time slot). It can be observed that, in symmetric configurations where each node transmits an equal number of bits per clock tick, the latency is proportional to the NCL. Consequently, minimizing the NCL is equivalent to minimizing latency in such settings.
Remark 2.
(Communication load of data placement) A worst-case analysis of the data placement phase reveals a transmission overhead of $O(\mathsf{N}^{2}r)$ to place $\mathsf{N}r$ files, given the $O(\mathsf{N})$ maximum path length in a ring network. This results in a normalized communication load of $O(\mathsf{N}r)$ per file. In many application scenarios, this cost would become negligible when the placement is amortized over multiple computation rounds. Furthermore, in scenarios where placement represents local data acquisition (e.g., sensing/satellite systems), the cost is application-dependent and remains beyond the scope of this work.
Remark 3.
(Relation to the MapReduce framework [9, 31]) The system with an arbitrary number of files and output functions can be reduced to the above setting through appropriate padding and batching of files and functions. In particular, our model can be interpreted within the MapReduce framework [9] by considering the special cases where each output function is computed either by all nodes (all-gather) or by exactly one node (all-to-all), under the additional constraint that each computing node has a limited broadcast distance. Specifically, consider the MapReduce-style system with $\mathsf{N}$ nodes are assigned $\mathsf{Q}\in\mathbb{N}^{+}$ output functions which takes input the $\mathsf{M}\in\mathbb{N}^{+}$ files. Let $g\triangleq\frac{\mathsf{M}}{\mathsf{N}},p\triangleq\frac{\mathsf{Q}}{\mathsf{N}}\in\mathbb{N}^{+}$, which can be satisfied by injecting empty files or functions into the system when they are not integers [9]. We can partition the $\mathsf{M}$ input files into $\mathsf{N}$ disjoint batches $\{\mathcal{B}_{1},\ldots,\mathcal{B}_{\mathsf{N}}\}$, each containing $g$ files. Let $\mathcal{W}_{k}\subseteq[\mathsf{Q}]$ be the index set of output functions assigned to node $n_{k}$. By viewing each batch $\mathcal{B}_{i}$ as a whole file $\hat{w}_{i}$, the MapReduce-style system can be transformed into our system model with $\mathsf{N}$ files and $\mathsf{N}$ nodes:
  • $\mathcal{W}_{i}=\mathcal{W}_{j}$ and $|\mathcal{W}_{i}|=Q$ for all $i,j\in[\mathsf{N}]$ corresponds to the all-gather computing, where each output function is computed by all nodes.
  • $\mathcal{W}_{i}\cap\mathcal{W}_{j}=\emptyset$ and $|\mathcal{W}_{i}|=|\mathcal{W}_{j}|=\frac{\mathsf{Q}}{\mathsf{N}}$ for all $i\neq j$ corresponds to the all-to-all computing, where each output function is computed by exactly one node.
In the MapReduce-style system [9], the computation load is defined identically to Definition 1. The normalized communication load is also defined analogously, differing from our definition only by an additional scaling factor. In most of the paper, we take $\mathsf{M}=\mathsf{Q}=\mathsf{N}$; this setting simplifies exposition while preserving the generality of the provided results and analytic techniques.
@@ II-B1 Reverse Carpooling
Our transmission strategy is built on reverse carpooling [34] that realizes the benefits of network coding. As demonstrated in Fig. 3 (a), where nodes $n_{1}$ and $n_{3}$ want to exchange packets through a relay node $n_{2}$. After receiving packets $P_{a}$ and $P_{b}$ sent from two different nodes on the opposite sides, the relay $n_{2}$ broadcasts the linear combination $P_{1}\oplus P_{2}$. Each of $n_{1}$ and $n_{3}$ can then obtain the desired packet by subtracting the packet it sent previously from the mixed packet. Therefore, with network coding, it needs $3$ transmissions. In contrast, the traditional forwarding approach requires $4$ transmissions (two for each packet). In general, reverse carpooling involves two information flows that traverse a common path in opposite directions. For example, as shown in Fig. 3(b), the path $(n_{1},n_{2},n_{3},n_{4})$ is shared by two flows, one from $n_{1}$ to $n_{4}$ and the other from $n_{4}$ to $n_{1}$. Without reverse carpooling, each hop must forward packets for the two flows separately, resulting in double transmissions. In contrast, reverse carpooling utilizes XOR-based broadcasting for bidirectional traffic. This allows a single transmission on shared links to serve two flows simultaneously. Consequently, under a pipelined transmission schedule, the number of required transmissions on the shared path is reduced by half, leading to an approximately $50\%$ transmission saving. There will be multiple information flows in our ring network. The reverse carpooling technique can benefit the transmission when some broadcasts form the structure as Fig. 3(a).
@@
Figure 3: (a) Reverse carpooling with 33 nodes. (b) Reverse carpooling for two flows. $P_{a}^{(t)}$ and $P_{b}^{(t)}$ are the packets sent by nodes $n_{1}$ and $n_{4}$ at time clock $t$, respectively. At clock $t$, node $n_{2}$ broadcasts $P_{a}^{(t-1)}\oplus P_{b}^{(t-2)}$, and node $n_{3}$ broadcasts $P_{a}^{(t-2)}\oplus P_{b}^{(t-1)}$. It effectively enables the two flows to traverse a common path without interfering with each other.
@@ II-B2 Efficient Broadcasting Using Network Coding
    if $k>1$, each $\alpha_{i}\in A$ transmits the sum of the two information symbols it received in phase $2$, round k1k-1.
@@
Following Algorithm 1, at the end of round $k$, each node in $A$ and $B$ can receive two new symbols from the two source nodes that are $2k$ and 2k12k-1 nodes away along the circle, respectively. The algorithm will finish after at most $\frac{\mathsf{N}}{4}$ rounds, and all nodes can decode each input symbol. We can conclude that the NCL is
@@ III Main Results
For the all-gather computing system with a computation load $r\in\{1,\ldots,\mathsf{N}\}$ and a broadcast distance $d\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$ over the considered ring networks with $\mathsf{N}$ nodes, the following NCL is achievable
@@
$\mathsf{T}_{1}^{\star}(r,d)\geq\frac{\mathsf{N}-r}{2d},$ (7)
for any file placement. Moreover, we have
@@
Whenever $2d\mid(\mathsf{N}-r)$, the proposed scheme is exactly optimal.
@@
The lower convex of all points $\big\{\big(r,\left\lceil\frac{\mathsf{N}-r}{2d}\right\rceil\big):r\in\{1,\ldots,\mathsf{N}\}\big\}$ is achievable for general $1\leq r\leq\mathsf{N}$ by using memory-sharing.
@@
For the all-to-all computing over the ring-based networks, we obtain the achievable NCL in Theorem 2 and a lower bound on NCL under the cyclic placement in Theorem 3.
@@
$\mathsf{T}_{\textnormal{2-cyc}}^{\textnormal{ach}}(r,d)=\left\{\hskip-5.0pt\begin{array}[]{cc}\vskip 5.0pt\frac{d}{2}\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor^{2}+\frac{d}{2}\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor+\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor\left(\left\lceil(\mathsf{N}-r)/2\right\rceil\right)_{d}+\left\lceil(\mathsf{N}-r)/2\right\rceil,&\!2r-1\leq\!d\!,\\ \vskip 5.0pt\frac{d}{2}\!\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-r+1}{d}\right\rfloor^{2}\!\!+\!\frac{3d}{2}\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-r+1}{d}\right\rfloor\!\!+\!r\!-\!1\!+\!\left(\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-r+1}{d}\right\rfloor\!\!+\!\!2\right)\left(\lceil\frac{\mathsf{N}-r}{2}\rceil\!-\!d\right)_{d},&\!\!r\!-\!1\!<\!d\!\leq\!2(r\!-\!1),\\ \frac{d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor^{2}\!+\!\frac{d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor\!+\!\left(\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor\!+\!1\right)\left(\left\lceil(\mathsf{N}-r)/2\right\rceil\!-\!d\right)_{d},&\!d\!\leq r\!-\!1,\end{array}\right$. (12)
where $(x)_{d}$ denotes $x\bmod d$ for $x\in\mathbb{Z}$, and the dominant term of $\mathsf{T}_{\textnormal{2-cyc}}^{\textnormal{ach}}(r,d)$ is $\frac{\mathsf{N}^{2}}{8d}$ for different ranges of $d$ and $r$.
@@
For the all-to-all computing system with a given computation load $r\in\{1,\ldots,\mathsf{N}-1\}$ and a broadcast distance $d\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$ over the considered ring networks with $\mathsf{N}$ nodes, a lower bound for the NCL under the cyclic placement is given as:
$\mathsf{T}_{\textnormal{2-cyc}}^{\star}(r,d)\geq\max_{s\in\{1,\ldots\mathsf{N}\}}\frac{s(\mathsf{N}-s-r+1)}{2d},$ (13)
where the right-hand side is maximized at $s=\lceil\frac{\mathsf{N}-r+1}{2}\rceil$, giving the scaling of lower bound as $\frac{(\mathsf{N}-r+1)^{2}}{8d}$.
@@
To expose the asymptotic scaling more transparently (ignoring integer rounding), the achievable NCL in Theorem 2 can also be expressed as
$\mathsf{T}_{\textnormal{2-cyc}}^{\textnormal{ach}}(r,d)=\left\{\begin{array}[]{cc}\vskip 5.0ptO\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{3(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8d}\right),&\!2r-1\leq\!d,\!\\ \vskip 5.0pt\!\!\!O\!\!\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!3r)\!+\!\frac{3\mathsf{N}-5r}{4}\!+\!\frac{9r^{2}+4(\mathsf{N}-r+1)}{8d}\right),&\!\!r\!-\!1\!<\!d\!\leq\!2(r\!-\!1),\\ O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8d}\right),&\!d\!\leq r\!-\!1.\end{array}\right$. (17)
It is readily observed that the achievable NCL in Theorem 2 are order-optimal. Specifically, for $\mathsf{N}\gg r$, the multiplicative gap is bounded as $\frac{\mathsf{T}_{\textnormal{2-cyc}}^{\textnormal{ach}}(r,d)}{\mathsf{T}_{\textnormal{2-cyc}}^{\star}(r,d)}\leq\frac{\mathsf{N}^{2}}{(\mathsf{N}-r+1)^{2}}+\epsilon$, where $\epsilon\to 0$ as $\mathsf{N}\to\infty$. This ratio asymptotically approaches $1$ as $\mathsf{N}\to\infty$. Next, we further present a converse bound for arbitrary file placement when $r\leq\frac{\mathsf{N}}{2}$ and $d\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$.
Theorem 4.
For the all-to-all computing system with a given computation load $r\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$ and a broadcast distance $d\in\{1,\ldots,\lfloor\frac{\mathsf{N}}{2}\rfloor\}$ over the considered ring networks with $\mathsf{N}$ nodes, a lower bound for the NCL is given as:
$\mathsf{T}_{\textnormal{2}}^{\star}(r,d)\geq\max_{s\in\{1,\ldots\mathsf{N}\}}\frac{s(\mathsf{N}-sr)}{2d},$ (18)
where the right-hand side is maximized at $s=\lceil\frac{\mathsf{N}}{2r}\rceil$ (or $\lfloor\frac{\mathsf{N}}{2r}\rfloor$), giving the scaling of lower bound as $\frac{\mathsf{N}^{2}}{8dr}$.
Proof.
Please see the converse proof in Appendix B. ∎
The performance gap between the achievable NCL in Theorem 2 and the converse bound for arbitrary placement in Theorem 4 is characterized $\frac{\mathsf{T}_{\textnormal{2-cyc}}^{\textnormal{ach}}(r,d)}{\mathsf{T}_{\textnormal{2}}^{\star}(r,d)}\leq\frac{(\mathsf{N}-r)^{2}}{\mathsf{N}^{2}/r}+\epsilon$, where $\epsilon\to 0$ as $\mathsf{N}\to\infty$. For small values of $r$, this ratio asymptotically approaches $r$. This gap is expected since Theorem 4 provides a general lower bound over all placements, while Theorem 2 is obtained under the cyclic placement. Our results indicate that the cyclic placement is particularly competitive when the computation load is relatively small, which is reasonable in practical distributed computing applications. A figure illustrating this comparison is provided in Fig. 5.
Figure 5: Comparison of achievable NCL and lower bound of all-to-all. The achievable NCL is obtained under the cyclic placement, while the lower bound is derived for arbitrary placement.
Figure 6: Comparison of achievable NCL and lower bound of all-gather.
Figure 7: Comparison of achievable NCL and lower bound of all-to-all under cyclic file placement.
When the computation load is relatively large $r\in\{\lceil\frac{\mathsf{N}}{2}\rceil,\ldots,\mathsf{N}\}$, without constraints on data placement, an optimal result is obtained when $d=1$ as shown in the following theorem.